OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (4, -4) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

do you hhave any idea..??

OpenStudy (anonymous):

r and theta can be found at looking at the graph :)

OpenStudy (anonymous):

where's the graph..?

OpenStudy (anonymous):

M giving answer...just a moment..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got (6,pi/4) but im confused cuz

OpenStudy (anonymous):

these are the answers (4 square root of 2, 45°), (-4 square root of 2, 225°) (4 square root of 2, 135°), (-4 square root of 2, 315°) (4 square root of 2, 225°), (-4 square root of 2, 45°) (4 square root of 2, 315°), (-4 square root of 2, 135°)

OpenStudy (anonymous):

i converted it into polar

OpenStudy (anonymous):

1) Let A be the point (x, y) = (4, - 4) => x = 4 and y = - 4 r = √(x² + y²) = √(16 + 16) = √32 = ± 4 √2 tan Θ = -4/4 = - 1 => Θ = (i) 315º or - 45º ; (ii) 135º or - 225 Hence, the Polar Coordinates are (i) (4√2, 315º) (ii) (-4 √2, 135º)

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i dont understand the polar coordinate part

OpenStudy (anonymous):

how you get the digits 315 and 135

OpenStudy (anonymous):

look the values of tan=-1 at 135 and 315 only..you have to take that values at which tan gives -1 in between 0 to 360 so these are two at which tan gives -1..

OpenStudy (anonymous):

ahhh ok

OpenStudy (anonymous):

you get it fully..?

OpenStudy (anonymous):

is it possible to find polar coordinates from a polar coordinate?

OpenStudy (anonymous):

for example (6,31^degrees)

OpenStudy (anonymous):

wht means.??

OpenStudy (anonymous):

how would i find the all it's polar coordinates from looking at (6,31^degrees)